# Efficient Long-Range Machine Learning Force Fields for Liquid and Materials Properties

### v1

@@ 3.1 Organic models
We train and test three models specific to tasks for organic molecules. OrganicMPNICE is a multitask model trained to the SPICE dataset (B97M-D3BJ/def2-TZVPPD),41, 42 OrbNet Denali dataset (B97X-D3/def2-TZVP),43 and a proprietary dataset of 11M finite structures at the B97X-D3BJ/def2-TZVPD level. OrganicMPNICETB is a delta-learned10 model trained to correct GFN2-xTB44 energies for the same dataset. OrganicCrystalsMPNICE is a multitask model trained to the same dataset, alongside a set of 400K molecular crystals evaluated using PBE-D3. For this model we report errors using two output heads, where it is of interest. The output head being currently used is indicated in parenthesis, for example, OrganicCrystalsMPNICE (PBE-D3) indicates the use of the PBE-D3 output head of this multi-task model. Unless otherwise specified the B97X-D3BJ/def2-TZVPD output head is used. More detailed information on how these datasets were constructed as well as tables summarizing the datasets and models trained can be seen in sections 5.1 and 5.2.
@@ 4 Conclusions and Future Developments
For use in predicting relative energetics of conformations of finite organic molecules, as well as isomerization and tautomerization, we suggest the use of OrganicMPNICETB, with OrganicMPNICE a close second and can be used in cases where efficiency is paramount or for which periodic simulations are required. For molecular dynamics simulations of organic liquids, we suggest the use of OrganicMPNICE. For simulations of molecular crystals, including ranking for crystal structure prediction, we suggest the use of OrganicCrystalsMPNICE (PBE-D3). For bulk inorganic materials and inorganic surfaces, we suggest the use of InorganicMPNICE. For approximation of the structure of organometallic complexes, we suggest using HybridMPNICE or HybridMPNICEO.
@@
While these results are notable, there are some areas which require further development. In particular, MPNICE cannot rigorously distinguish systems with distinct spin states, and so is unable to predict quantities such as singlet-triplet gaps or magnetic ordering in materials. While HybridMPNICE and HybridMPNICEO have been demonstrated to achieve reasonable approximations of organometallic complexes, they still exhibit decreased performance versus domain-specific models, in particular failing to obtain liquid solvent densities within 10 of experimental values. Use of the presented pretrained models to unseen classes of material - such as reactive datasets, surfaces, and defects - are extrapolative in nature and thus subject to proportional skepticism, and fine-tuning is likely to be necessary for greater accuracy in some applications. As we have not yet tested performance for interfaces between organic liquids and materials, this limits confidence in zero-shot simulations of such interfaces, which are of significant importance in materials science. There is still much room to develop techniques to train a general MLFF which can effectively produce a single PES for systems with disparate reference levels of theory.
@@ 5.1.1 Lifesciences Dataset
In order to cover some bonding patterns not common to druglike datasets we have enhanced the “dark matter universe” (DMU), developed along with the well known GDB-13 dataset.87 We utilize substitution rules, similar to the construction of GDB-1788 to expand the DMU to the element set stated above. This source of molecules is very large and contains many examples of highly reactive molecules. We selected all molecules with fewer than four non-hydrogen (heavy) atoms as well as 2500 additional, randomly selected molecules. Each of these molecules was optimized and examples were selected from 20 ps of MD, resulting in 135,000 examples of small molecules.
To cover common intermolecular interactions of drug-like molecules we utilize a sampling protocol which involves performing single component molecular dynamics of disordered systems, followed by extraction of clusters. Once a molecule is chosen we perform stereochemical enumeration and limited conformational sampling. This yields a structurally diverse set of stereoisomers that we pack into a disordered systems (liquid). This liquid is then optimized, followed by NVT MD at the optimized volume and 150 ps of NPT MD. Clusters are extracted from the NPT dynamics. Here, we select molecules in two ways. We first filter our element enhanced DMU dataset, using filters similar to Fink et al.87 to yield a set of relatively stable organic molecules. We then select the following sets of molecules: all systems with three or fewer heavy atoms, a random selection in each of a number of special groups we found to be error prone with our prior MLFF (halogen containing molecules, halogenated aromatic groups, hydrazines, nitro groups, cyano groups and sulfonamides), and finally, a random sample of the full filtered, element-enhanced DMU.
@@
DFT calculations for finite systems labeled with B97X-D3BJ/def2-TZVPD were performed with Psi489 v1.6 using the same settings as in ref. 90:
@@ 5.2 Model training
6 Code Availability
7 Data Availability
Reference data for all tests are publicly available at https://github.com/leifjacobson/MLFF_test_data. {acknowledgement} JLW would like to thank Steven Dajnowicz for providing the geometries for Pt/Ir complexes.

### v2

@@ 3.1 Organic models
We train and test three models specific to tasks for organic molecules. OrganicMPNICE is a multitask model trained to the SPICE dataset (B97M-D3BJ/def2-TZVPPD),41, 42 OrbNet Denali dataset (B97X-D3/def2-TZVP),43 and a proprietary dataset of 11M finite structures at the B97X-D3BJ/def2-TZVPD level. OrganicMPNICETB is a delta-learned10 model trained to correct GFN2-xTB44 energies for the same dataset. OrganicCrystalsMPNICE is a multitask model trained to the same dataset, alongside a set of 400K molecular crystals evaluated using PBE-D3. For this model we report errors using two output heads, where it is of interest. The output head being currently used is indicated in parenthesis, for example, OrganicCrystalsMPNICE (PBE-D3) indicates the use of the PBE-D3 output head of this multi-task model. Unless otherwise specified the B97X-D3BJ/def2-TZVPD output head is used. More detailed information on how these datasets were constructed as well as tables summarizing the datasets and models trained can be seen in sections 9.1 and 9.2.
@@ 4 Conclusions and Future Developments
For use in predicting relative energetics of conformations of finite organic molecules, as well as isomerization and tautomerization, we suggest the use of OrganicMPNICETB, with OrganicMPNICE a close second, which can be used in cases where efficiency is paramount, or when periodic simulations are required. For molecular dynamics simulations of organic liquids, we suggest the use of OrganicMPNICE. For simulations of molecular crystals, including energetic ranking of polymorphs for organic crystal structure prediction, we suggest the use of OrganicCrystalsMPNICE (PBE-D3). For bulk inorganic materials and inorganic surfaces, we suggest the use of InorganicMPNICE. For approximation of the structure of organometallic complexes, we suggest using HybridMPNICE or HybridMPNICEO.
We have not extensively explored the limit of MPNICE model capacity; it is therefore possible that training significantly larger models, as in the recent 150 million parameter UMA models,87 would result in less loss of accuracy when training hybrid models such as HybridMPNICE. However, doing so will necessarily increase inference cost, limiting large-scale MD simulations and thus the accessible condensed-phase properties. While a broadly trained MLFF can yield reasonable simulation results for, e.g., simple liquids (see Sec. 3.1.4), the diversity of applications for general pretrained MLFFs virtually guarantees the necessity of fine tuning, particularly in targeted discovery campaigns requiring highly accurate predictions in complex environments such as battery materials and biological molecules. For such use-cases the advantage of general pretrained models lies in the learned internal representations (as in section 3.1.6), which allows significantly more data efficient training and model stability as compared to training a model from scratch. The ease of fine tuning, and the efficiency of the resulting model, relies on the efficiency of the initial architecture and pretrained models. Rapid evolution of MLFF technology over the next 3-5 years will reveal what is necessary and sufficient for efficient, predictive simulations of complex biological and materials science systems. We believe that the MPNICE architecture and pretrained models represent a useful step in this direction.
@@
While the results presented here are notable, there are some areas which require further development. In particular, MPNICE cannot rigorously distinguish systems with distinct spin states, and so is unable to predict quantities such as singlet-triplet gaps or magnetic ordering in materials. While HybridMPNICE and HybridMPNICEO have been demonstrated to achieve reasonable approximations of organometallic complexes, they still exhibit decreased performance versus domain-specific models, in particular failing to obtain liquid solvent densities within 10 of experimental values. Use of the presented pretrained models to unseen classes of material - such as reactive datasets, surfaces, and defects - are extrapolative in nature and thus subject to proportional skepticism, and fine-tuning is likely to be necessary for greater accuracy in some applications. We expect this to be at least partially mediated in the future through incorporation of the rapidly evolving availability of additional training data, such as the recently released OMol25 dataset, which was reported after this work was completed and includes extensive organometallic and reactive sampling at the B97M-V level of theory.88 Interfaces between organic liquids and materials, for which the reference level of theory is not well defined, are additionally not tested here. This limits confidence in zero-shot simulations of such interfaces, which are of significant importance in materials science. There is still much room to develop techniques to train a general MLFF which can effectively produce a single PES for systems with disparate reference levels of theory.
5 Author Contributions
J.L.W. conceptualized the model, developed software, trained and validated the inorganic and hybrid models and authored the manuscript. R.D.G. analyzed results for the matbench discovery benchmark. G. A. performed a case study on LiAlO2. Y.W. performed fine tuning for EA prediction and performed DLPNO-CCSD(T) computations on TorsionNet500. A.A.F. developed software. X.X. curated water data, conceptualized the free energy algorithm and performed hydration free energy calculations. J.S. developed software, curated data and edited the manuscript. B.S. performed PBE-D3 computations on organic crystals. L.D.J. conceptualized the project, developed software, curated organic data, trained and validated the organic models and authored the manuscript. M.D.H. and R.A. reviewed and edited the manuscript. L.D.J., R.A.F., M.D.H., K.L. and R.A. supervised the project.
6 Conflicts of interest
The authors declare the following competing financial interest(s): R.A.F. has a significant financial stake in, is a consultant for, and is on the Scientific Advisory Board of Schrodinger, Inc.
7 Code Availability
8 Data Availability
Reference data for validation tests that are not already publicly available can be found at https://github.com/leifjacobson/MLFF_test_data. This includes reference energy data and geometries for the tautobase tautomer test, TorsionTest2000, DLPNO-CCSD(T) energies for TorsionNet500, the organic crystal test geometries for 20 crystal forming organic molecules with PBE-D3 energies as well as equilibrated initial geometries used for liquid density evaluation. {acknowledgement} J.L.W. would like to thank Steven Dajnowicz for providing the geometries for Pt/Ir complexes.
@@ 9.1.1 Lifesciences Dataset
In order to cover some bonding patterns not common to druglike datasets we have enhanced the “dark matter universe” (DMU), developed along with the well known GDB-13 dataset.89 We utilize substitution rules, similar to the construction of GDB-1790 to expand the DMU to the element set stated above. This source of molecules is very large and contains many examples of highly reactive molecules. We selected all molecules with fewer than four non-hydrogen (heavy) atoms as well as 2500 additional, randomly selected molecules. Each of these molecules was optimized and examples were selected from 20 ps of MD, resulting in 135,000 examples of small molecules.
To cover common intermolecular interactions of drug-like molecules we utilize a sampling protocol which involves performing single component molecular dynamics of disordered systems, followed by extraction of clusters. Once a molecule is chosen we perform stereochemical enumeration and limited conformational sampling. This yields a structurally diverse set of stereoisomers that we pack into a disordered systems (liquid). This liquid is then optimized, followed by NVT MD at the optimized volume and 150 ps of NPT MD. Clusters are extracted from the NPT dynamics. Here, we select molecules in two ways. We first filter our element enhanced DMU dataset, using filters similar to Fink et al.89 to yield a set of relatively stable organic molecules. We then select the following sets of molecules: all systems with three or fewer heavy atoms, a random selection in each of a number of special groups we found to be error prone with our prior MLFF (halogen containing molecules, halogenated aromatic groups, hydrazines, nitro groups, cyano groups and sulfonamides), and finally, a random sample of the full filtered, element-enhanced DMU.
@@
DFT calculations for finite systems labeled with B97X-D3BJ/def2-TZVPD were performed with Psi491 v1.6 using the same settings as in ref. 92: